# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 5 Discussion and Future Work
Table 9: Limitations of Existing Micronutrient Status Assessment Methods and Opportunities to Address Them
@@
Limited clinical relevance - Compare new approaches to the clinical gold-standard
@@
5.1 Clinically-Relevant Innovations
Generally, we find that current innovative approaches to biofluid analysis for micronutrient status assessment lack clinical relevance (Sections 3.2 to 3.5). To overcome this limitation and develop effective, meaningful assessment methods, future research could prioritize demonstrating clinical utility and shift focus toward clinical relevance as a primary goal. At a high level, this requires an interdisciplinary approach to technological research, and an understanding that novel, technological methods are best utilized as an aid to clinical expertise. Micronutrition is fundamentally a clinical field, so all future solutions must contextualize their work within a clinical application.
It is imperative to benchmark new approaches against established clinical standards. Such comparative validation is crucial to demonstrate the relevance and efficacy of a novel method, and to determine its potential as a viable substitute for the clinical gold standard. Only a limited number of the reviewed methodologies conduct such comparative analyses with state-of-the-art clinical practices. To aid in this endeavor, gold-standard methods for the assessment of each micronutrient are depicted in Tables 17, 18, and 19 in the Appendix. We note that ELISA tests are sufficient in most cases, as long as the standard protocol is followed.
Since this validation can be difficult and expensive (Section 3.1), there is potential for studies to apply new assessment methods among patient populations that are routinely assessed by clinicians. Often, the clinical gold standard is already being used to determine the status of these individuals. Thus, there is an opportunity to collaborate with medical professionals working in these populations, who may provide access to clinical test results for comparative analysis as well as valuable feedback. Selecting populations at risk for imbalance, such as candidates for bariatric surgery or individuals with diabetes, may be useful because they are frequently surveyed and assessed for nutritional status pre and post-treatment.
Researchers could also understand and integrate clinical literature about the relevance of different biofluid samples (Section 2.2), utilizing them where appropriate. For example, several reviewed papers claim to measure in-body micronutrient status from sweat, despite clinical evidence that sweat is an inaccurate representation of in-body status [14, 13]. Clinical relevance can be further improved by assuring novel methods can measure the same biomarkers used by clinicians across the spectrum of deficiency to excess.
5.2 Comprehensive Approaches through Individualized and Multi-Modal Solutions
Micronutrition is a broad and complex field requiring a holistic and individualized approach (Section 1). To develop effective and applicable technologies, it is essential to embrace precision nutrition assessment and multi-modal sensing. Precision nutrition leverages technology to account for individual differences in diet, metabolism, demographics, lifestyle, and more [124, 159, 90, 192]. This allows nutritional diagnoses and interventions to be fine-tuned to a specific patient. Future micronutrient assessment methods could integrate elements of precision nutrition by considering these individual differences. AI and ML will be powerful tools in this endeavour, allowing for the consideration of the varied aspects of micronutrition during an assessment.
Multi-modal sensing is a promising opportunity to incorporate precision nutrition into new assessment methods. Current methods conduct biofluid analysis (Section 3) or physiological sensing (Section 4) in isolation. New technologies can combine these approaches with conventional health data, such as physical activity, demographics, and other clinical information. A multi-modal approach that leverages both biofluid and physiological analyses can reveal more relevant insights into nutritional status, alleviate over-reliance on subjective assessments (like dietary reports and physical exams; Section 4.1), and ultimately create a comprehensive overview of an individual’s health.
Multi-modal sensing also allows for the analysis of multiple micronutrients simultaneously, a functional necessity since micronutrient imbalances rarely manifest alone [12]. Current biofluid analysis methods often focus on assessing a single micronutrient, but precision nutrition emphasizes the need to consider all aspects of an individual’s health to gather the most relevant insights. Micronutrients interact in complex ways (Tables LABEL:tab:_char-ws, LABEL:tab:_char-fs, LABEL:tab:_char-min), and there is opportunity for more methods that target several micronutrients concurrently to understand their comprehensive effects on individual health. If it is not feasible for a single device to measure multiple micronutrients, future work could combine several devices or measurement modes reviewed here together in a unified solution.
We urge calls by Brown et al. [21] for more micronutrition data, as progress towards precision assessment can be accelerated by access to data on individual nutrition prognoses (Section 3.5). Innovative micronutrient status assessment devices and other health sensors could be utilized during clinical studies and interventions to collect the necessary data for assessing micronutrient status in a personalized manner. Additionally, studies could analyze micronutrient status changes within and between individuals over time, leading to important insights and discoveries. Useful data for research and analysis includes patient demographic information, previous clinical history, and all nutritional screening or assessments conducted through the course of treatment. Images of physical symptoms, biochemical analysis results, and data from a wearable device such as a FitBit could be critical to valuable insights. Increased data availability could enable clinicians to gain greater insights and discoveries, allow for applying data-hungry AI and ML models [36], and pave the way for early detection of micronutrient imbalance. Diversity across cultures, gender, race, and other demographic and socioeconomic factors is essential to the global value of new research in precision micronutrient assessment. The impact of such factors should be carefully analyzed and reported, creating a more well-rounded understanding of nutrition that minimizes bias. To achieve this, accessible PoC devices will be needed to collect data at scale, while AI and ML tools will be useful to mine for insights.
5.3 Accessible and Noninvasive Point-of-Care Devices
This work highlights the critical need for accessible and noninvasive methods to assess micronutrient status. Accessible assessment can both aid in data generation, through POoC sample analysis [90], and be a byproduct of it, through a greater general understanding of micronutrition. Our review of existing methods reveals that current technologies have yet to fully address these limitations. The solution to this issue involves leveraging commodity hardware, open-source designs, prioritizing ease-of-use for non-experts, and reducing invasiveness in assessment techniques.
As technology becomes increasingly available to the general public, they can be used to power accessible micronutrient status assessment methods. Interfacing with commodity smart devices as analyzers or additional data sources will continue to be valuable for future work. Smartphones and smartwatches have powerful built-in sensing capabilities with the potential to reveal physiological micronutrition insights. Their utility can be expanded by adding accessories like multiplexed electrochemical sensor chips or microfluidic pumps to perform micronutrient status analyses on biofluids.
Improving ease of use for non-clinicians could significantly enhance accessibility, allowing these devices to be used at home or in the field. Methods that allow users to gain important insights into their micronutritional status and understand the impact on their overall health should be both understandable and actionable for the general user, without needing to consult a medical professional. Although this level of accessibility is a future goal, achieving it would provide individuals with the most direct way to obtain micronutritional status and related insights to take control of their health.
To be more accessible, future methods should be minimally invasive. Current methods strongly rely on invasively-collected biofluid samples, such as serum and plasma. One promising approach is to utilize alternative biofluids, such as urine, which is less invasive to obtain and clinically relevant for assessing micronutrient status (Section 2.2). An even less invasive strategy is to bypass biofluid samples altogether, leveraging on-body devices such as wearables. For example, some research has shown promise in developing on-body spectroscopic devices [35, 182], and others have demonstrated the simultaneous quantification of multiple micronutrients from the spectra of a complex sample using novel ML approaches [112, 113]. Accessible, noninvasive approaches could allow for continuous monitoring of micronutrient status, providing valuable insights and offering a novel approach to micronutrient assessment.
6 Conclusion
@@
This article provides a comprehensive review of accessible and noninvasive in-body micronutrient status assessment methods. We take a critical look at a wide variety of research in the field, based on analyses of both biofluids (Section 3) and physiology (Section 4). These current methods, such as assays, electrochemistry, spectroscopy, optical sensors, and AI & ML, are evaluated in the context of their performance and overall clinical relevance. Several tables are contributed throughout the paper and in the Appendix, intended to gather the most relevant information about micronutrients and their assessment together into an intuitive reference. Our major contributions to micronutrient status assessment as a research area include (1) general background information on micronutrients for a non-clinical audience, (2) synthesis of existing technological and clinical micronutrient status assessment methods, split into biofluid and physiological-based techniques, (3) recommendations for future directions to develop accessible and noninvasive assessment methods, and (4) a unique focus on clinical relevance throughout.
The review synthesized existing micronutrient status assessment methods, split into analyses based on biofluids and physiology. We find that biofluid-based assessments benefit from established biomarkers of micronutrient status and more attention on micronutrients in general, but suffer from highly specific methodologies that frequently rely on blood matrices and lack a comparison to a reliable clinical standard of in-body assessment. Assessments that consider physiology must combat the associative nature of physiological symptoms, the bias inherent in diet self-reporting, as well as the sparseness of micronutrient-specific clinical research or technological advancements. By taking a unique, clinical perspective, this review reveals that no technological solution attempts to consider the holistic nature of micronutrition. Where the clinical standard is to consider both biofluid and physiological analyses together, the focus of current research is predominantly on the application of either biofluid or physiological analysis for the assessment of one single micronutrient at a time.
This review concludes with a discussion of the limitations and gaps in current micronutrient status assessment methods, highlighting actionable opportunities for future research. Three specific opportunities are outlined that require immediate attention to expedite progress in this field. As summarized in Table 9, these opportunities address gaps related to limited clinical relevance, the absence of holistic and comprehensive approaches, and the high invasiveness and inaccessibility of current methods. These challenges can be overcome through clinically relevant innovations, comprehensive individualized multi-modal solutions, and the development of accessible and non-invasive PoC devices. Advancing towards these goals would enable individuals to effectively combat the ’hidden hunger’ of micronutrient malnutrition using accessible, non-invasive methods.
@@ Appendix A Tables
Table 10: Alphabetized List of Abbreviations in Main Text
@@
BMVA breast milk vitamin A NIR near infrared
BP blood pressure PBS phosphate buffered saline
Bphen bathophenanthroline PNS parasympathetic nervous system
BRI body roundness index PoC point of care
CBC complete blood count PPG photoplethysmogram
CoV coefficient of variance PR pulse rate
CRP C-reactive protein PRV pulse rate variability
CUN-BAE Clinical University of Navarra body adiposity estimator PWV peak wavelength value
CVD cardiovascular disease RBP retinol binding protein
ECG electrocardiogram RDA Recommended Dietary Allowance
EDA electrodermal activity RF radio frequency
EDR estimated daily requirement RMSE root mean square error
ELISA enzyme-linked immunosorbent assay ROI region of interest
GSH glutathione SNS sympathetic nervous system
HF high-frequency SPR surface plasmon resonance
@@
Table 11: Characteristics of Micronutrients: Water-Soluble Vitamins. Information from [12, 34, 115, 17, 119]
@@
Vitamin B2 (riboflavin); 2 coenxyme derivatives (FMN and FAD) Critical to energy metabolism, cell development and functionality, and metabolism of fats, drugs, and steroids (maintains homocysteine levels) Small amounts in liver, heart, kidneys Lack of evidence Vegitarian athletes, pregnant and lactating women and their infants, people who are vegan and/or consume little milk, people with riboflavin transporter deficiency Absorption decreased by copper, zinc, iron, manganese intake; deficiency associated with those of folate, pyridoxine, nacin Alcoholism, Chronic intestinal failure None
@@
Vitamin B7 (biotin) Critical to the metabolism of proteins, fats, and carbohydrates into energy Most stored in liver None Those with biotinidase deficiency, chronic alcohol exposure, and pregnant and breastfeeding women Alcoholism, chronic intestinal failure Anticonvulsants
@@
Vitamin B12 (cobalamin) [5] Critical to CNS development and functionality, RBC formulation, DNA synthesis, conversion of homocysteine 80% in liver; 1-5 mg (thousands times more than daily consumption); can last 2-5 years, 1-3 by some sources None Women, elderly, black people, those with low socioeconomic status, who have had gastrointestinal surgery, are vegitarian/vegan Absorption decreased by excess vitamin C Alcoholism, chronic intestinal failure, chronic (atrophic) gastritis, Liver diseases, obesity post bariatric surgery, critical illness Gastric acid inhibitors, metformin
Vitamin C (ascorbic acid) Required in synthesis of collagen and neurotransmitters, used in protein metabolism, and critical to immune function High concentrations in cells and tissues, WBC, eyes, adrenal glands, pituitary gland, and brain; total content 300 mg (near acute deficiency) to 2g Yes (mild nausea, diarrhea, cramps) Smokers, those with low food variety, any disease causing oxidative stress Shown to regenerate other antioxidants (ex vitamin E) Alcoholism, chronic (atrophic) gastritis, obesity post bariatric surgery, critical illnesss Chemotherapy/radiation, 3-hydroxy-3-methylglutaryl coenzyme A reductase inhibitors
Table 12: Characteristics of Micronutrients: Fat-Soluble Vitamins. Information from [12, 34, 115, 17, 119]
@@
Vitamin A Critical for vision, cell growth, immune and reporductive functions Most in liver (about 6 months), some in eyes Yes Infants, pregnant women in low/middle income/developing countries Absorption decreased by zinc deficiency Alcoholism, chronic intestinal failure, Inflammatory bowel diseases, liver diseases, obesity post bariatric surgery, cystic fibrosis Orlistat, retinoids (results in toxicity)
Vitamin D; 2 forms: 25(OH)D (calcidiol) and 1,25(OH)D (calcitriol) [144] Bone groth and strength, absorption and control of calcium, reducing inflammation Fatty tissue and liver Yes Breastfed infants, adults 20-39, those with kidney/liver dysfunction, with dark skin, limited sun exposure, conditions limiting fat absorption Magnesium is critical to activation and binding, function is heavily interwound with Calcium Alcoholism, chronic intestinal failure, chronic (atrophic) gastritis, Inflammatory bowel diseases, liver diseases, obesity post bariatric surgery, chronic renal failure, critical illness Orlistat, statins, steroids, thiazide diuretics
@@
Table 13: Characteristics of Micronutrients: Minerals. Information from [12, 34, 115, 17, 119]
@@
Iron [58, 55] Essential to oxygen transport through hemoglobin, energy metabolism, physical growth, neurological development, cell functioning, and hormone synthesis 60% in blood hemoglobin, rest as ferritin in liver, spleen, bone marrow, muscles Yes (especially those with hemochromatosis and elderly) Infants, young children, teen girls, those who are pregnant (especially if Mexican-American or black), premenopausal, in food-insecure households, have increased menstaral bleeding Absorption increased by Vitamin C intake, Absorption decreased by zinc, calcium, manganese intake and copper deficiency Chronic intestinal failure, chronic (atrophic) gastritis, inflammatory bowel diseases, obesity post bariatric surgery, critical illness, cancer, heart failure Levodopa, levothyroxine, proton pump inhibitors
Copper Cofactor in energy production, iron absorption, neuropeptide activation, and synthesis of connective tissue and neurotransmitters 50-120mg total, 95% carried by ceruloplasmin; 2-3 months in skeleton and muscle; tightly regulated, only  1mg/d loss in bile Yes Those who are pregnant Absorption decreased by high zinc Chronic intestinal failure, obesity post bariatric surgery, chronic renal failure, critical illness, celiac disease, menkes disease
Zinc [130] Physical growth and development, cellular metabolism, and immune functions 85% in skeletal muscle and bone,  0.1% in plasma where 70% of that is bound to albumin; 1.5g women, 2.5g men total Yes Children, teens, exclusively breastfed infants, those who are pregnant, vegitarian/vegan, have eating disorders, malabsorption, gastrointestinal disorders Absorption decreased by high calcium/iron Alcoholism, chronic intestinal failure, inflammatory bowel diseases, liver diseases, obesity post bariatric surgery, chronic renal failure, critical illness, sickle cell disease, HIV Antibiotics, penicillamine, piuretics
Iodine Thyroid gland function, protein synthesis, metabolism and enzyme activity 70-80% in thyroid gland; 15-20 mg total Yes Infants, those who are pregnant, use uniodized salt, are in regions with iodine-deficient soils Absorption decreased by iron intake and selenium deficiency Anti-thyroids, angiotensin-converting enzyme inhibitors, potassium-paring diuretics
@@
Magnesium Regulates several chemical reactions, including blood glucose and blood pressure regulation, DNA, RNA, and protein synthesis, proper muscle and nerve functioning, bone development, and calcium and potassium ion transport Approx 25g; 50-60% in bone, <1% in serum (tightly controlled), rest in soft tissue Yes Elderly Absorption increased by Vitamin D Alcoholism, gastrointestinal disease, bariatric surgery, T2D Bisphosphonates, antibiotics, diuretics, proton pump inhibitors
Table 14: Physiological Symptoms of Micronutrient Deficiencies: Water-Soluble Vitamins. Information from [123, 37, 39, 140, 142, 115, 17, 119]
@@
Vitamin B2 Conjunctiva inflamation/grittines (angular blepharitis), Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/iritation, swollen/sticky eyelid, photophobia Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), dry/swollen/ulcerated lips (cheilosis), redness in lips and tongue, swollen/inflammed/smooth tongue (glossitis), Atrophied papillae Correlated with fatigue
Vitamin B3 Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/iritation, swollen/sticky eyelid Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), dry/swollen/ulcerated lips (cheilosis), redness in lips and tongue, swollen/inflammed/smooth tongue (glossitis), Atrophied papillae, inflamed gums (gingivitis) Pellagra Correlated with fatigue Biomarkers indicate insufficiency far before clinical symptoms appear
@@
Vitamin B6 Conjunctiva inflamation/grittines (angular blepharitis), conjuntiva pallor, Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/iritation, swollen/sticky eyelid Excessive thinness, hapalonychia Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), swollen/inflammed/smooth tongue (glossitis), dry/swollen/ulcerated lips (cheilosis), Atrophied papillae, redness in lips and tongue Anaemia, cardiomyopathies/heart failure Supplementation improves blood pressure, reported to help regulate SNS Correlated with fatigue Borderline and mild status may not present symptoms for months or years; Radler and Lister [140] say "deficiency often occurs within 2 months of inadequacy"

### v2

@@ 5.1. Evaluating Emerging Methods
Table 11. Limitations of Existing Micronutrient Status Assessment Methods and Opportunities to Address Them.
@@
Limited clinical relevance - Compare new approaches to the clinical optimal reference
@@
5.2. Future Opportunities to Address Limitations of Micronutrient Assessment
5.2.1. Clinically Relevant Innovations
We find that current innovative approaches to biofluid analysis for micronutrient status assessment lack clinical relevance (Sections 3.2 to 3.5). To overcome this limitation, future research could shift focus toward clinical relevance as a primary goal. This requires an interdisciplinary approach and recognition that technology is most effective when it complements clinical expertise, especially as micronutrition is inherently a clinical field where all solutions must be grounded in practical applications. Clinical, practical grounding can also be done by benchmarking new approaches against established optimal reference standards, detailed in Table LABEL:tab:_assessment. Comparative validation is essential to demonstrate a novel method’s relevance and its potential as a viable substitute for the clinical optimal reference standard. We note that appropriate use of ELISA tests are sufficient in most cases.
We also find it is important for researchers to prioritize clinically relevant biofluids and ensure measurements reflect meaningful biomarker levels. For instance, reviewed papers claim to leverage sweat, yet sweat is an unreliable indicator of micronutrient status (Baker and Wolfe, 2020; Baker, 2019). Clinical relevance improves when methods align with the biomarkers used across the full spectrum of deficiency to excess. However, future research can also validate which biosamples and biomarkers may be suitable replacements for the optimal reference standard for specific micronutrients, and in specific usage contexts. Since clinical validation can be challenging (Section 3.1), studies could apply new assessment methods in patient populations already monitored by clinicians, offering opportunities for collaboration and access to clinical reference data for validation. Targeting at-risk groups, like bariatric surgery candidates or individuals with diabetes, is especially useful as they are routinely assessed for nutritional status.
5.2.2. Comprehensive Approaches through Individualized and Multi-Modal Solutions
Micronutrition is complex and requires a holistic, individualized approach (Section 1). Advancing assessment technologies calls for precision nutrition and multi-modal sensing that account for differences in diet, metabolism, and lifestyle (of Public Health, 2020; Srinivasan et al., 2017; Kirk et al., 2021; Zmora and Elinav, 2021). While current methods often analyze biofluids (Section 3) or physiology (Section 4) separately, combining them with clinical data, like treatment history and anthropomorphic measurements, can uncover patterns and offer a more complete view of micronutritional status. Multi-modal sensing also enables simultaneous analysis of multiple micronutrients, essential due to their complex interactions (Bailey et al., 2015). (Tables LABEL:tab:_char-ws, LABEL:tab:_char-fs, LABEL:tab:_char-min). When single-device measurement is not feasible, future solutions could integrate multiple tools into a unified system.
We support calls for more micronutrition data (Brown et al., 2021), as access to individual nutrition profiles can accelerate progress toward precision assessment (Section 3.5). Innovative assessment devices and health sensors could be used in clinical studies to collect personalized micronutrient data over time. Valuable information includes demographics, clinical history, nutritional assessments, symptom images, biochemical results, and wearable data. Greater data availability would support deeper clinical investigations and aid early detection of imbalances. Disease etiology lies at the intersection of comprehensive clinical assessment, preventative care, and clinical relevance. Etiology stresses the multifactorial (and deeply individual, social, and cultural) nature of disease progression (White, 2020). Comprehensive assessments, and data thereof, can contribute to the etiological understanding of micronutrient malnutrition and ultimately drive precise prevention strategies.
AI and ML tools can help integrate these factors, uncovering latent patterns across a vast quantity and variety of micronutrition data (Côté and Lamarche, 2022). We have discussed how micronutritional assessments are made more powerful when they are conducted and compared repeatedly over time. Accessible, multi-modal sensing would enable routine collection of health data relevant to micronutritional status, which can be combined and compared in real time, and en masse, by ML models. However, the computational and environmental costs of AI and ML models, especially large-scale generative AI models, must also be considered. It will also be essential to ensure demographic and cultural diversity to produce relevant and equitable models. Scalable and sustainable data collection through accessible PoC devices could be essential to achieving this.
5.2.3. Accessible and Noninvasive Point-of-Care Devices
This work highlights the critical need for accessible and noninvasive methods to assess micronutrient status. Such methods can generate valuable data though PoC assessments (Kirk et al., 2021) and therefore improve the overall understanding of micronutrition. Current technologies fall short in accessibility, but potential solutions lie in leveraging commodity hardware, open-sourcing designs, prioritizing ease of use for non-experts, and utilizing less invasive biosamples. Smartphones and smartwatches, with built-in sensors, offer promising platforms for physiological insights. Their capabilities can be further expanded with features or accessories like electrochemical sensor chips or microfluidic pumps for biofluid analysis. As a result, multi-modal solutions with broad sensing and analysis capabilities will become more broadly available and low-cost.
Methods could be built considering some existing wearable devices. The Empatica E4 and Polar H10 offer high-fidelity HRV and autonomic data, which is associated vitamin B12 and iron status (Section 2.3) (Schuurmans et al., 2020; Giles et al., 2016). Oura Ring, WHOOP, Fitbit, and Garmin devices track sleep, heart rate, and recovery metrics, which have been linked to nutrient-related fatigue and dysfunction (Henriksen et al., 2022; Miller et al., 2020; Benedetto et al., 2018; Kolla et al., 2020). Additionally, Dexcom continuous glucose monitors, while focused on glucose, demonstrate the clinical viability of wearable biochemical sensing and may inform future designs for micronutrient monitoring (Garg et al., 2022). While these tools vary in accuracy, especially across sensor types and populations, they highlight growing opportunities for accessible physiological monitoring relevant to micronutrition. Minimally invasive alternatives, such as urine testing (Section 2.2) instead of serum or plasma, can enhance accessibility. Even more promising are wearable, on-body sensing devices which use spectroscopy and ML to monitor micronutrients continuously without the need for biosamples (Crocombe, 2018; Watson et al., 2022; Monakhova et al., 2010; Monakhova and Mushtakova, 2017). When care is taken to design innovative on-body, PoC assessments that are open-source and easy to use by laypeople (in addition to being accessible and noninvasive), comprehensive micronutrient status assessment can become viable for resource limited communities as well as individuals at home.
6. Conclusion
@@
This article provides a comprehensive review of accessible and noninvasive methods for assessing in-body micronutrient status, focusing on biofluids (Section 3) and physiological (Section 4) approaches. We evlauate current techniques such as assays, electrochemistry, spectroscopy, optical sensors, and AI/ML for performance and clinical relevance. Key contributions include: (1) background on micronutrients for non-clinical audiences, (2) a synthesis of biofluid- and physiology-based methods, (3) future directions for noninvasive and accessible assessment, and (4) a unique focus on clinical applicability. Several summary tables provide an intuitive reference throughout.
This review synthesizes micronutrient status assessment methods based on biofluid and physiological analyses. Biofluid approaches benefit from established biomarkers but often rely on blood samples and lack clinical validation. Physiological assessments face challenges like weak symptom associations, self-report bias, and limited micronutrient-specific research. From a clinical perspective, no current technology holistically integrates both types of analysis, despite this being standard in practice. Most research focuses on assessing a single micronutrient using either biofluid or physiological data alone. This review outlines three key opportunities to advance micronutrient status assessment: improving clinical relevance, adopting holistic multi-modal approaches, and reducing invasiveness to enhance accessibility (Table 11). Addressing these gaps through innovative, non-invasive, and individualized PoC solutions could empower individuals to better manage micronutrient malnutrition.
Acknowledgments
Special thanks to Sofia Luis for her assistance editing the revised version of this paper, and to the anonymous reviewers for their constructive feedback.
@@ Appendix A Tables
Table 12. Alphabetized List of Abbreviations in Main Text
@@
BP blood pressure NIR near infrared
Bphen bathophenanthroline PBS phosphate buffered saline
BRI body roundness index PNS parasympathetic nervous system
CBC complete blood count PoC point of care
CoV coefficient of variance PPG photoplethysmogram
CRP C-reactive protein PR pulse rate
CUN-BAE Clinical University of Navarra body adiposity estimator PRV pulse rate variability
CVD cardiovascular disease PWV peak wavelength value
ECG electrocardiogram RBP retinol binding protein
EDA electrodermal activity RDA Recommended Dietary Allowance
EDR estimated daily requirement RF radio frequency
ELISA enzyme-linked immunosorbent assay RMSE root mean square error
GSH glutathione ROI region of interest
HF high-frequency SNS sympathetic nervous system
HMVA human milk vitamin A SPR surface plasmon resonance
@@
Table 13. Characteristics of Micronutrients: Water-Soluble Vitamins. Information from (Bailey et al., 2015; Collier, 2010; Mueller, 2017; Berger et al., 2022; National Institutes of Health, datea)
@@
Vitamin B2 (riboflavin); 2 coenzyme derivatives (FMN and FAD) Critical to energy metabolism, cell development and functionality, and metabolism of fats, drugs, and steroids (maintains homocysteine levels) Small amounts in liver, heart, kidneys Lack of evidence Vegetarian athletes, pregnant and lactating people and their infants, people who are vegan and/or consume little milk, people with riboflavin transporter deficiency Absorption decreased by copper, zinc, iron, manganese intake; deficiency associated with those of folate, pyridoxine, niacin Alcoholism, Chronic intestinal failure None
@@
Vitamin B7 (biotin) Critical to the metabolism of proteins, fats, and carbohydrates into energy Most stored in liver None Those with biotinidase deficiency, chronic alcohol exposure, and pregnant and breastfeeding people Alcoholism, chronic intestinal failure Anticonvulsants
@@
Vitamin B12 (cobalamin) (Ada’s Medical Knowledge Team, 2023) Critical to CNS development and functionality, RBC formulation, DNA synthesis, conversion of homocysteine 80% in liver; 1-5 mg (thousands times more than daily consumption); can last 2-5 years, 1-3 by some sources None Women, elderly, black people, those with low socioeconomic status, who have had gastrointestinal surgery, are vegetarian/vegan Absorption decreased by excess vitamin C Alcoholism, chronic intestinal failure, chronic (atrophic) gastritis, Liver diseases, obesity post bariatric surgery, critical illness Gastric acid inhibitors, metformin
Vitamin C (ascorbic acid) Required in synthesis of collagen and neurotransmitters, used in protein metabolism, and critical to immune function High concentrations in cells and tissues, WBC, eyes, adrenal glands, pituitary gland, and brain; total content 300 mg (near acute deficiency) to 2g Yes (mild nausea, diarrhea, cramps) Smokers, those with low food variety, any disease causing oxidative stress Shown to regenerate other antioxidants (ex vitamin E) Alcoholism, chronic (atrophic) gastritis, obesity post bariatric surgery, critical illness Chemotherapy/radiation, 3-hydroxy-3-methylglutaryl coenzyme A reductase inhibitors
Table 14. Characteristics of Micronutrients: Fat-Soluble Vitamins. Information from (Bailey et al., 2015; Collier, 2010; Mueller, 2017; Berger et al., 2022; National Institutes of Health, datea)
@@
Vitamin A Critical for vision, cell growth, immune and reproductive functions Most in liver (about 6 months), some in eyes Yes Infants, pregnant people in low/middle income/developing countries Absorption decreased by zinc deficiency Alcoholism, chronic intestinal failure, Inflammatory bowel diseases, liver diseases, obesity post bariatric surgery, cystic fibrosis Orlistat, retinoids (results in toxicity)
Vitamin D; 2 forms: 25(OH)D (calcidiol) and 1,25(OH)D (calcitriol) (Ross and Institute of Medicine (2011), U. S.) Bone growth and strength, absorption and control of calcium, reducing inflammation Fatty tissue and liver Yes Breastfed infants, adults 20-39, those with kidney/liver dysfunction, with dark skin, limited sun exposure, conditions limiting fat absorption Magnesium is critical to activation and binding, function is heavily interwoven with Calcium Alcoholism, chronic intestinal failure, chronic (atrophic) gastritis, Inflammatory bowel diseases, liver diseases, obesity post bariatric surgery, chronic renal failure, critical illness Orlistat, statins, steroids, thiazide diuretics
@@
Table 15. Characteristics of Micronutrients: Minerals. Information from (Bailey et al., 2015; Collier, 2010; Mueller, 2017; Berger et al., 2022; National Institutes of Health, datea)
@@
Iron (Gloria, date; Gerber, 2024) Essential to oxygen transport through hemoglobin, energy metabolism, physical growth, neurological development, cell functioning, and hormone synthesis 60% in blood hemoglobin, rest as ferritin in liver, spleen, bone marrow, muscles Yes (especially those with hemochromatosis and elderly) Infants, young children, teen females, pregnant people (especially if Mexican-American or Black), premenopausal, in food-insecure households, have increased menstrual bleeding Absorption increased by Vitamin C intake, Absorption decreased by zinc, calcium, manganese intake and copper deficiency Chronic intestinal failure, chronic (atrophic) gastritis, inflammatory bowel diseases, obesity post bariatric surgery, critical illness, cancer, heart failure Levodopa, levothyroxine, proton pump inhibitors
Copper Cofactor in energy production, iron absorption, neuropeptide activation, and synthesis of connective tissue and neurotransmitters 50-120mg total, 95% carried by ceruloplasmin; 2-3 months in skeleton and muscle; tightly regulated, only  1mg/d loss in bile Yes Pregnant people Absorption decreased by high zinc Chronic intestinal failure, obesity post bariatric surgery, chronic renal failure, critical illness, celiac disease, menkes disease
Zinc (Pat Bass, 2023) Physical growth and development, cellular metabolism, and immune functions 85% in skeletal muscle and bone,  0.1% in plasma where 70% of that is bound to albumin; 1.5g females, 2.5g males total Yes Children, teens, exclusively breastfed infants, pregnant people, vegetarian/vegan, have eating disorders, malabsorption, gastrointestinal disorders Absorption decreased by high calcium/iron Alcoholism, chronic intestinal failure, inflammatory bowel diseases, liver diseases, obesity post bariatric surgery, chronic renal failure, critical illness, sickle cell disease, HIV Antibiotics, penicillamine, piuretics
Iodine Thyroid gland function, protein synthesis, metabolism and enzyme activity 70-80% in thyroid gland; 15-20 mg total Yes Infants, pregnant people, use uniodized salt, are in regions with iodine-deficient soils Absorption decreased by iron intake and selenium deficiency Anti-thyroids, angiotensin-converting enzyme inhibitors, potassium-paring diuretics
@@
Magnesium Regulates several chemical reactions, including blood glucose and blood pressure regulation, DNA, RNA, and protein synthesis, proper muscle and nerve functioning, bone development, and calcium and potassium ion transport Approx 25g; 50-60% in bone, ¡1% in serum (tightly controlled), rest in soft tissue Yes Elderly Absorption increased by Vitamin D Alcoholism, gastrointestinal disease, bariatric surgery, T2D Bisphosphonates, antibiotics, diuretics, proton pump inhibitors
Table 16. Physiological Symptoms of Micronutrient Deficiencies: Water-Soluble Vitamins. Information from (of Nutrition and Dietetics, 2017; De and Seshadri, 2012; DiBaise and Tarleton, 2019; Radler and Lister, 2013; Reber et al., 2019; Mueller, 2017; Berger et al., 2022; National Institutes of Health, datea)
@@
Vitamin B2 Conjunctiva inflammation/grittiness (angular blepharitis), Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/irritation, swollen/sticky eyelid, photophobia Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), dry/swollen/ulcerated lips (cheilosis), redness in lips and tongue, swollen/inflamed/smooth tongue (glossitis), Atrophied papillae Correlated with fatigue
Vitamin B3 Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/irritation, swollen/sticky eyelid Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), dry/swollen/ulcerated lips (cheilosis), redness in lips and tongue, swollen/inflamed/smooth tongue (glossitis), Atrophied papillae, inflamed gums (gingivitis) Pellagra Correlated with fatigue Biomarkers indicate insufficiency far before clinical symptoms appear
@@
Vitamin B6 Conjunctiva inflammation/grittiness (angular blepharitis), conjuntiva pallor, Redness/fissures in eyelid corners (Angular Palpebritis), conjunctiva redness/irritation, swollen/sticky eyelid Excessive thinness, hapalonychia Bilateral cracks/redness at corners of lips/mouth (angular cheilosis), swollen/inflamed/smooth tongue (glossitis), dry/swollen/ulcerated lips (cheilosis), Atrophied papillae, redness in lips and tongue Anemia, cardiomyopathies/heart failure Supplementation improves blood pressure, reported to help regulate SNS Correlated with fatigue Borderline and mild status may not present symptoms for months or years; Radler and Lister (2013) say ”deficiency often occurs within 2 months of inadequacy”